### v1

# ProcessPainter: Learn Painting Process from Sequence Data

Abstract.
The painting process of artists is inherently stepwise and varies significantly among different painters and styles. Generating detailed, step-by-step painting processes is essential for art education and research, yet remains largely underexplored. Traditional stroke-based rendering methods break down images into sequences of brushstrokes, yet they fall short of replicating the authentic processes of artists, with limitations confined to basic brushstroke modifications. Text-to-image models utilizing diffusion processes generate images through iterative denoising, also diverge substantially from artists’ painting process. To address these challenges, we introduce ProcessPainter, a text-to-video model that is initially pre-trained on synthetic data and subsequently fine-tuned with a select set of artists’ painting sequences using the LoRA model. This approach successfully generates painting processes from text prompts for the first time. Furthermore, we introduce an Artwork Replication Network capable of accepting arbitrary-frame input, which facilitates the controlled generation of painting processes, decomposing images into painting sequences, and completing semi-finished artworks. This paper offers new perspectives and tools for advancing art education and image generation technology.
Diffusion model, Image generation, Painting process
Figure 1. This paper accomplishes the generation of a painting process that mimics human artists, transitioning from abstract to specific. Each row in the figure represents a different application: (1) generating painting processes from text prompt (Text2Painting), (2) converting artworks into painting processes (Image2Painting), (3) completing semi-finished paintings (Semi2Complete).
[…]
1. Introduction
The step-by-step painting process has a wide range of applications, making it highly suitable for educational, entertainment, and professional fields. Generating a painting sequence that replicates the style of human artists is a long-desired but challenging goal(Hertzmann, 2003). Previous studies have employed stroke-based rendering methods using stepwise greedy search(Vanderhaeghe and Collomosse, 2012), recurrent neural networks(Aksan et al., 2020; Cao et al., 2019), reinforcement learning(Schaldenbrand and Oh, 2021), and parameter searching(Tong et al., 2022; Hu et al., 2023) to simulate the human painting process. The Intelli-Paint(Singh et al., 2022) utilizes semantic segmentation and hierarchical rendering to mimic the sequence of human painting. However, these methods have limited success in emulating the diverse painting processes of different artists, subjects, and art forms. Moreover, the process of human painting involves rich strategies and intelligence; painting subjects and backgrounds separately is just one of many strategies. These methods do not learn from human painting process key-frames and merely focus on minimizing similarity loss with input images by preventing brush strokes.
In the era of generative AI, the intersection of artificial intelligence and creative arts has seen significant advances in image synthesis technology, allowing the conversion of textual descriptions into visual representations(Zhou et al., 2021; Yang et al., 2023; Cho et al., 2024). Models typically based on diffusion(Li et al., 2024; Ge et al., 2023) processes have opened new horizons for artistic exploration, yet they still fail to generate step-by-step painting processes akin to those of human artists, making them unsuitable for creating step-by-step painting tutorials. Because painting is generally a gradual instantiation process, moving from abstract to specific, from macro to detail(Liu et al., 2021), which is far removed from how diffusion models generate images through a denoising process. Decoding the latent from the denoising process through a VAE decoder only yields blurry images, not meaningful painting processes.
To address this, we introduce ProcessPainter, a data-driven solution that can generate vivid and realistic painting processes with controllable styles. ProcessPainter leverages temporal models trained on synthetic datasets and artists’ painting sequences, defining the task of generating painting processes as a specific video generation problem. However, training or fine-tuning video models require massive amounts of data, and high-quality serialized painting video data are scarce. Videos documenting painting processes on the internet often include obstructions from the painter’s hands, changes in camera angles, and zoom, making them unsuitable as training data. To innovate, we propose using an SBR-based method to construct synthetic datasets for training foundational painting models, enabling the generation of painting processes from textual descriptions. We then fine-tune the model with a small number of human painting sequences using LoRA(Hu et al., 2021) to learn the styles and process strategies of human artists. Painting LoRA not only replicates the artists’ final visual styles but also their unique creative processes, providing a detailed observation of ’how’ art is made, a feature largely missing in traditional diffusion models and SBR applications.
In painting education and practice, students copying classic works and teachers completing and modifying students’ semi-finished pieces are two important teaching steps. To enable ProcessPainter as an AI tool to assist in these steps, we propose the Artwork Replication Network, which implements two practical functionalities: converting artworks to process key-frames and completing semi-finished paintings. Specifically, we trained an Artwork Replication Network that can accept any frame as a reference condition. During inference, inputting the reference image as the last frame enables generating a sequence from a blank canvas to the reference image. If the semi-finished reference image is input as a mid-sequence condition, ProcessPainter can progressively complete the reference image. We tested our method on various real-world images and paintings of different styles, and our experimental results demonstrate that our method can generate vivid painting processes with a high degree of anthropomorphism and artistic sense. Furthermore, our method is plug-and-play; by replacing different style Unets or mounting LoRA models, our approach can generate painting processes in various styles, such as oil painting, sketching, and ink painting.
The contributions of our paper are summarized as follows:
  • We introduce a novel method for generating painting processes from textual descriptions, conceptualizing it as a video generation process.
  • We propose a method that involves pre-training the painting model on synthetic data, followed by training Painting LoRA to learn specific strategies and styles from a few human painting key-frames.
  • We propose an Artwork Replication Network that accommodates control through arbitrary-frame images, enabling the conversion of reference images into painting process and the completion of semi-finished paintings.
2. Related Works
2.1. Stroke-based Rendering
The problem of teaching machines “how to paint” has been extensively studied in the context of stroke-based rendering (SBR), which focuses on the recreation of non-photorealistic imagery through appropriate positioning and selection of discrete elements such as paint strokes or stipples(Hertzmann, 2003). Early SBR approaches involved either a greedy, iterative search or required user interaction to address the decomposition problem (Haeberli, 1990; Litwinowicz, 1997; Hertzmann, 2022). In recent years, recurrent neural networks (RNNs) and reinforcement learning (RL) have been extensively employed to generate strokes sequentially (Ha and Eck, 2017; Zhou et al., 2018; Xie et al., 2013; Singh et al., 2022). Moreover, integrating adversarial training (Nakano, 2019) has emerged as a viable method for generating non-deterministic stroke sequences. adversarial updates of actors, critics, and discriminators. Stylized Neural Painting(Zou et al., 2021; Kotovenko et al., 2021) proposed stroke optimization strategy that iteratively searches optimal parameters for each stroke and and is possible to be optimized jointly with neural style transfer. Although the aforementioned methods achieve the painting process from abstract to specific, these methods still significantly differ from the creative processes of human artists. This is because the painting processes of different artists and subjects vary significantly, making it challenging for a single method to simulate all. Recently, the emergence of Stable Diffusion(Rombach et al., 2022) technology has significantly improved the effectiveness of generative images. Exploring the generation of painting processes based on diffusion models is a promising new field. To this end, we propose ProcessPainter, which fine-tunes diffusion models using data from artists’ painting processes to learn their true distributions.
2.2. Text2image Diffusion Model
Recent studies have demonstrated that diffusion models are capable of generating high-quality synthetic images, effectively balancing diversity and fidelity. Models based on diffusion models or their variants, such as those documented in (Nichol et al., 2021; Ramesh et al., 2022; Saharia et al., 2022), have successfully addressed the challenges associated with text-conditioned image synthesis. Stable Diffusion (Rombach et al., 2022), a model based on the Latent Diffusion Model (Rombach et al., 2022), incorporates text conditioning within a UNet framework to facilitate text-based image generation, establishing itself as a mainstream model in image generation (Podell et al., 2023; Li et al., 2023; Feng et al., 2023). Fine-tuning pre-trained image generation models can enhance their adaptation to specific application scenarios, as seen in techniques like LoRA (Hu et al., 2021) and DreamBooth (Ruiz et al., 2023). For theme control in text-to-image generation, several works (Kumari et al., 2023; Ye et al., 2023; Wang et al., 2024a) focus on custom generation for defined pictorial concepts, with ControlNet (Zhang et al., 2023a) additionally offering control over other modalities such as depth information. These Diffusion models have revolutionized the field of image generation with their powerful generative capabilities. However, the iterative denoising process is completely different from human painting, and cannot be mimicked with a brush, thus offering limited guidance for painting instruction. This paper attempts to use pre-trained diffusion models to generate painting processes that resemble those of artists.
Figure 2. Overall schematics of our method. During the training phase, we first pre-train the Painting model and Artwork Replication Network on 40,000 synthetic data points. Then, we fine-tune the Painting LoRA model on a small amount of artists’ painting process data. During inference, ProcessPainter generates the painting process step by step from a reference image, producing the final painting as the last frame. It can also refine a partially completed image based on textual descriptions and the input image. When no reference image is provided, ProcessPainter generates the painting process solely from textual descriptions.
2.3. Video Generation Models
The foundational frameworks for image and video generation share similarities, primarily encompassing three pathways: Generative Adversarial Networks (GANs) (Tian et al., 2021; Shen et al., 2023), autoregressive models (Le Moing et al., 2021; Ge et al., 2022; Hong et al., 2022), and diffusion models. Within this context, stable diffusion-based video generation models utilize training of 3D U-Nets to restore high-fidelity videos from Gaussian noises (Guo et al., 2023; Lin and Yang, 2024; Blattmann et al., 2023; Khachatryan et al., 2023). To enhance the controllability of generated videos, recent advancements in motion customization video generation models employ multimodal control signals to influence the visual appearance and cinematographic effects of the outputs (Dai et al., 2023; Ma et al., 2024b, a). Specifically, MotionCtrl (Wang et al., 2023) facilitates precise control over camera and object motion, while ControlVideo inherits the architecture and weights from ControlNet (Zhang et al., 2023a), allowing for controllable video generation through motion sequences. Notably, there are no existing models for controllable video generation aimed at teaching painting processes. Thus, we propose ProcessPainter, which explores the creation of videos in various styles of painting processes, introducing a novel domain in stable diffusion-based video generation technology.
3. Method
In this section, we begin by exploring the preliminaries on diffusion models as detailed in Section 3.1. Then introduce the Network architecture of our method inSection 3.2, followed by detailed descriptions of the key modules: the Painting Model, Painting LoRA, and Artwork Replication Network. In Section 3.3, we discuss the construction methods of the dataset and conclude with the settings for the training and inference phases in Section 3.4.
3.1. Preliminary
Diffusion Model (DM). DM (Ho et al., 2020; Wang et al., 2024b) belongs to the category of generative models that denoise from a Gaussian prior $x_{T}$ to target data distribution $x_{0}$ by means of an iterative denoising procedure. The common loss used in DM is:
(1) $L_{\text{simple}}(\theta):=\mathbb{E}_{x_{0},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t)\|^{2}_{2}\right],$
where $x_{t}$ is a noisy image constructed by adding noise $\epsilon\sim\mathcal{N}(0,1)$ to the natural image $x_{0}$ and the network $\epsilon_{\theta}(\cdot)$ is trained to predict the added noise. At inference time, data samples can be generated from Gaussian noise $\epsilon\sim\mathcal{N}(0,1)$ using the predicted noise $\epsilon_{\theta}(x_{t},t)$ at each timestep $t$ with samplers like DDPM (Ho et al., 2020) or DDIM (Song et al., 2020).
Latent Diffusion Model (LDM). LDM (Rombach et al., 2022) is proposed to model image representations in the autoencoder’s latent space. LDM significantly speeds up the sampling process and facilitates text-to-image generation by incorporating additional text conditions. The LDM loss is:
(2) $L_{\text{LDM}}(\theta):=\mathbb{E}_{z_{0},t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(z_{t},t,\tau_{\theta}(c_{t}))\|^{2}_{2}\right],$
where $z_{0}$ represents image latents and $\tau_{\theta}(\cdot)$ refers to the BERT text encoder (Devlin et al., 2018) used to encode text description $c_{t}$.
Stable Diffusion (SD). SD is a widely adopted text-to-image diffusion model based on LDM. Compared to LDM, SD is trained on the large LAION (Schuhmann et al., 2021) dataset and replaces BERT with the pre-trained CLIP (Radford et al., 2021) text encoder.
3.2. Network Architecture
3.2.1. Overall Architecture.
The overall architecture of ProcessPainter is depicted in Fig. 2. The model is segmented into two primary components: a pre-trained stable diffusion model(Rombach et al., 2022) with Temporal-Attention(Guo et al., 2023), and an Artwork Replication Network. The parameters of the pre-trained stable diffusion, including the VAE and UNet, are fixed. We add a temporal module to each layer of the UNet, which is used to learn the inter-frame correlations of the progress key-frames. Additionally, an Artwork Replication Network accepts input from the image context. The Artwork Replication Network is integrated with the denoising UNet in an additive fashion, thereby offering the flexibility for seamless model switching.
3.2.2. Temporal-Painting Model and Painting LoRA
The key to generating painting sequences is that the entire sequence represents the transformation of the same image from abstract to detailed, ensuring consistency and correlation in content and composition across frames. To achieve this, we introduced the temporal attention module from AnimateDiff(Guo et al., 2023) into the UNet. This module, positioned after each diffusion layer, employs an inter-frame self-attention mechanism to assimilate information across different frames, ensuring smooth transitions and continuity throughout the sequence. This empirically validated training strategy grounds the model’s capability in inter-frame self-attention, where dividing the reshaped feature map along the temporal axis results in vector sequences of length $f$ (i.e., $\{z_{1},\ldots,z_{f};z_{i}\in\mathbb{R}^{(b\times h\times w)\times c}\}$), processed through self-attention blocks, i.e.
[…]
where $Q=W^{Q}z$, $K=W^{K}z$, and $V=W^{V}z$ are separate linear projections. This attention mechanism allows the current frame’s generation to utilize information from other frames.
Unlike video generation, the painting process undergoes more drastic changes from start to finish. The initial frame often consists of low-completion color blocks or sketches, while the final frame is a complete artwork. This presents a challenge for model training. To address this, we first pre-train the temporal module on a large synthetic dataset to learn the step-by-step painting processes using various Stroke- based-rendering methods. Then, we fine-tune the Painting LoRA model with 10-50 painting sequences from real artists. Notably, to decouple the artistic style of static works from the painting process strategies, we adopt a staged training approach for Painting LoRA. Initially, we train the Unet Spatial-Attention LoRA model using only the final frames of the painting sequences. Subsequently, we train the Temporal-Attention LoRA model on the full painting sequences. LoRA adds pairs of rank-decomposition matrices and optimizes only these newly introduced weights. By limiting the trainable parameters and keeping the original weights frozen, LoRA is less likely to cause catastrophic forgettingConcretely, the rank-decomposition matrices serve as the residual of the pre-trained model weights $\mathcal{W}\in\mathbb{R}^{m\times n}$.
(4) $\mathcal{W}^{\prime}=\mathcal{W}+\Delta\mathcal{W}=\mathcal{W}+AB^{T}$
where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{n\times r}$ are a pair of rank-decomposition matrices, $r$ is a hyper-parameter, which is referred to as the rank of LoRA layers. In practice, LoRA is only applied to attention layers, further reducing the cost and storage for model fine-tuning.
3.2.3. Artwork Replication Network.
In the previous section, we explored generating painting processes from textual descriptions. In this section, we introduce the Artwork Replication Network, which accepts arbitrary-frame inputs to achieve controllable painting process generation. It offers two practical functionalities: converting reference images to process key-frames and completing semi-finished paintings.
Similar to previous controllable generation methods(Zhang et al., 2023b, a; Nagahara, 2023), we introduce a variant of ControlNet to ensure that a specific frame $\tau$ in the generated sequence matches the reference image. As shown in Figure 2, during training and inference, the reference image is first encoded by the VAE encoder to obtain the latent representation $l_{i}$. We initialize a standard Gaussian noise tensor as the input for the Artwork Replication Network and replace the standard Gaussian noise at the position corresponding to frame $\tau$ with $l_{i}$. We add a Temporal-Attention layer after each Spatial-Attention layer in ControlNet to learn how the reference image should influence the entire sequence. The final output of the n-th layer of the denoising UNet $\hat{S_{n}^{u}}$ can be define as:
(5) $\hat{S_{n}^{u}}=S_{n}^{u}+\lambda ARN_{n}$
where $S_{n}^{u}$ is the output feature of the denoising UNet. $\lambda ARN_{n}$ is the output feature of the Artwork Replication Network. $n$ is the $n$-th layer in the Unet model. $\lambda$ is the coefficient for the degree of influence of the Artwork Replication Network.
During inference, The Artwork Replication Network supports multiple creative functions through different settings. When we set $\tau$ to the final frame, the reference image input will serve as the reconstruction target for the last frame of the painting process, allowing us to convert reference images to process key-frames. Similarly, when we set $\tau$ to the initial frames and input a partially completed painting as the reference, we achieve the effect of completing the reference image. The extent to which the network completes the partially finished painting depends on the setting of $\tau$. The closer $\tau$ is to the beginning, the greater the degree of painting completion.
Figure 3. Text to painting processe generation results. ProcessPainter can learn different process style from synthetic data.
3.3. Dataset Construction Method
Our dataset construction includes 30,000 synthetic painting sequences for pretraining, each with 8 frames at a resolution of 512x512, and 95 painting sequences from artists for LoRA trainting.
Synthetic Data. We first selected 10,000 images from the DiffusionDB dataset based on aesthetic scores, then converted the static images into painting sequences using three SBR methods: Learn-to-Paint, Stylized Neural Painting, and Paint Transformer. Since the last few hundred steps of SBR methods mainly optimize details with minimal changes to the overall image, we used non-uniform sampling to select 8 frames based on each method’s characteristics as the training set. Additionally, we proposed a method to simulate artists painting sub-regions from the foreground to the background for creating a synthetic dataset. Using Segment Anything(Kirillov et al., 2023) and Depth Anything(Yang et al., 2024), we segmented all objects in the image and added them to a blank canvas in order from nearest to farthest based on depth.
Artists’ Data. Besides synthetic data, we also collected a total of 95 painting sequences from three artists. The painting types include impasto portraits, landscape sketches, and coloring line art. For each artist’s work, we trained a separate Painting LoRA to learn their painting process strategies and styles. We add trigger word ”sks” in the image captions to describe the painting process we wish to train on.
3.4. Model Training and Inference
Figure 4. A Painting LoRA can be fine-tuned only on 10-50 sequences of artists’ painting process, which can effectively capture the characteristics of the artists’ painting process and the style of the final results.
Figure 5. Compare with stroke based rendering methods, our method provides a more precise reconstruction of the original images, and the painting process more closely resembles that of human artists. For different types of paintings, the strategies of the painting process can be controlled by switching the Painting Model or Painting LoRA.
(a) Converting artworks to painting process.
(b) Completing semi-finished paintings.
Figure 6. Given a reference image, ProcessPainter can generate a painting process that ensures specific frames match the reference image exactly.
In the training phase, we first pre-trained the Painting Model on a synthetic dataset. Then, we froze the parameters of the Painting Model and trained the Artwork Replication Network. For fine-tuning the Painting LoRA, we initially fine-tuned the Spatial-Attention LoRA using only the final frame to prevent unfinished painting works from damaging the image quality of the model. After this step, we froze the parameters of the Spatial-Attention LoRA and fine-tuned the Temporal-Attention LoRA using the complete painting sequence. We chose Prodigy as optimizer, which optimize the algorithm’s traversal of the optimization landscape by dynamically adjusting the learning rate for each parameter based on their past gradients, can be especially beneficial for LoRA training.
During the inference phase, when generating painting sequences from text, we do not use the Artwork Replication Network. For tasks involving image-to-painting sequence conversion and painting completion, we use the Artwork Replication Network to receive the reference input for specific frames. To ensure that the frames in the generated painting sequence that should match the input image do so as closely as possible, we employ the DDIM Inversion technique to obtain the initial noise of the reference image and replace the initial noise in the UNet for the specific frames.
4. Experiment
4.1. Experiment Setting
We conducted a pre-training session using 30,000 video sequences over 50,000 steps, enabling the painting model to effectively capture the transition from textual semantics to painting processes. Subsequently, the model underwent 50,000 steps of formal training, employing a randomly sampled Artwork Replication Network. Our frame sampling strategy included extracting between 0 and 3 frames, with a sampling probability of one-third for both the first and last frames and a normal distribution for the middle frames, peaking at the center. This strategy was designed to mimic realistic usage. We optimized the frame extraction method to facilitate convergence and to align with actual probability distributions. The learning rate was set at $2\times 10^{-5}$, with a batch size of 1, at a resolution of 512*512 pixels, and the number of frames per sequence fixed at 8. Pre-training and LoRA training were conducted on a NVIDIA A100 GPU. During inference, we employed DDIM sampler and DDIM inversion techniques for noise replacement, running each for 50 steps.
4.2. Text to Painting Process
In this section, we demonstrate the capability of a foundational model trained with synthetic data to generate paintings from text descriptions. As shown in Figure 3, the first row illustrates the painting process from rough to detailed using semi-transparent Bezier curves as brush strokes. The second row displays the step-by-step creation of an oil painting on a blank canvas using oil paint brushes. The third row shows the painting process from the top left corner to the bottom right of the canvas. These painting patterns were learned from different painting process strategies synthesized from data. Figure 4 displays the results of training the Painting LoRA model with a small number of key frames from human artists’ processes, showcasing three implementations: color block landscape sketching, line drawing coloring, and impasto portraits. The experiments demonstrated that with just a dozen or so samples and based on pre-trained painting model, we can replicate the strategies and styles of painters.
4.3. Image to Painting Process
In this section, we present the experimental results based on the Artwork Replication Network. As shown in Figure 6(a), by controlling the last frame with a reference image, we have achieved the conversion of artworks into painting processes. In this task, a text prompt is not necessary. We can control different strategies of the painting process by swapping out the painting model or the painting LoRA model, for instance, using oil paint brush strokes to render Van Gogh’s Starry Night or translucent strokes for Chinese ink paintings. When using a semi-finished piece as the reference image to control the first frame, we implemented text-guided painting completion capabilities. As displayed in Figure 6(b), the results from ProcessPainter align well with the text prompts, demonstrating good consistency. Based on the same semi-finished image, different results can be obtained through various text prompts.
4.4. Comparison and evaluation
Since our method is the first to achieve text-to-painting-process generation and step-by-step painting completion, there are no existing methods for comparison. In this section, we compare our method with state-of-the-art baseline methods in the Image-to-Painting-Process task. We chose 4 stroke-based-rendering methods for Comparison, they are LearnToPaint(Huang et al., 2019), Paint Transformer(Liu et al., 2021), Intelli-paint(Singh et al., 2022), and Stylized Neural Painting(Zou et al., 2021).
4.4.1. Qualitative Evaluation
As shown in Figure 5, although baseline methods can effectively replicate brushstroke processes, they are fundamentally designed to minimize the difference between the real image and the current canvas, resulting in a painting process that does not conform to human painting habits. Our method is based on stable diffusion, possesses a robust image prior, and has been fine-tuned with the painting processes of artists, thus aligning our results with the typical artist painting process. This approach showcases starting with broad color blocks and gradually adding details. Furthermore, the results from ProcessPainter exhibit more refined details and higher reconstruction accuracy relative to the reference image.
Table 1. Quantitative Evaluation of Reconstruction Consistency
Method Mean MSE $\downarrow$ Mean LPIPS $\downarrow$ Mean L1 $\downarrow$
LeranToPaint(Huang et al., 2019) 0.016181 0.033240 0.087082
Paint Transformer(Liu et al., 2021) 0.087695 0.153372 0.187685
Intelli-Paint(Singh et al., 2022) 0.247486 0.397536 0.350746
Stylized Neural Painting(Zou et al., 2021) 0.084447 0.141126 0.185832
Ours 0.014820 0.024517 0.082165
4.4.2. Quantitative Evaluation
In this section, we present the quantitative evaluation results. We evaluate the consistency between the output of the last frame in the image2painting task and the reference image. Specifically, we calculate the Mean Squared Error (MSE), Learned Perceptual Image Patch Similarity (LPIPS), and Least Absolute Deviations (L1). As a baseline comparison, we use LeranTopaint, Paint Transformer, Stylize neural painting, and Semantic RL, with results obtained from their official codebases and recommended parameters. Table 1 shows that our method achieves the highest reconstruction consistency compared with baseline methods.
4.5. Ablation Study
Figure 7. Ablation study on the consistency of the final frame reconstruction.
In this section, we conduct ablation studies to discuss the key designs in ProcessPainter that maintain consistency with the reference image. As shown in Figure 7, when both the Artwork Replication Network and the noise replacement strategy are employed, the results are highly consistent with the reference image. When only the Artwork Replication Network is used to maintain consistency, the results have limited consistency with the reference image. When only the noise replacement strategy is used, the consistency with the reference image is poor.
Table 2. User Preferences for Human-Like Painting Styles Comparison
Comparison Method Anthropomorphic (%) Preference (%)
Ours vs. LearnToPaint(Huang et al., 2019) 78.2 68.2
Ours vs. Paint Transformer(Liu et al., 2021) 80.4 65.9
Ours vs. Intelli-Paint(Singh et al., 2022) 84.5 78.4
Ours vs. Stylized Neural Painting(Zou et al., 2021) 78.6 71.6
4.6. User Study
The user study was conducted with 44 participants, each shown 28 paired painting sequences comparing our method to previous works. For each pair, participants were asked to select the painting sequence that best resembles the human painting style and which they preferred. Each sequence was presented as a GIF image with a total duration of 5 seconds. The results of the user study, shown in Table 2, suggest that our outcomes more closely resemble the painting processes of human artists and are more favored.
5. Limitation
The training sequences of paintings in this paper use 8 frames with a resolution of 512x512. Due to GPU memory constraints, it is difficult to generate sequences with more frames or higher resolution. Another limitation is that, although many digital painting tools can export the creation process of artists, obtaining the painting processes of human artists, compared to static images, remains challenging.
6. Conclusion
In this paper, we introduce ProcessPainter, a novel framework that bridges the gap between digital image synthesis and traditional artistic practices by generating detailed painting processes from text descriptions. By training temporal models on both synthetic datasets and artists’ painting sequence, ProcessPainter not only replicates the final visual styles of artists but also their unique creative processes. This approach provides a granular view of the artistic creation ’how-to’ that is often missing in traditional diffusion model applications. We proposed the Artwork Replication Network, which enables controllable painting process generation, allowing for the conversion of reference images into process key-frames and the completion of semi-finished paintings. Our contributions to the field are substantial, providing a new methodology for art creation, analysis, and education that promises to enrich both the understanding and practice of digital and traditional arts.